A social psychologist wants to test if extroversion is significantly related to a person’s general health. He gets extroversion scores and a measure of general health from the participants’ medical providers. The data are listed below. Higher health extroversion scores mean greater extroversion, while high health scores = good health.

Complete the appropriate statistical test and hypothesis testing using the 4 step approach. State your conclusion in APA format. Alpha level = .05.

Participant

extrov X

health Y

X squared Y squared (X) (Y)
A

101

33 3333
B 93 40 3720
C 85 24 2040
D 111 46 5106
E 104 30 3120
F 89 28 2492
Total
Answer #1

r (Using Excel function CORREL(x,y)) = 0.6997

n = 6, Degrees of freedom: df = 6-2 = 4, Level of significance = 0.05

H0: ρ = 0, Extroversion is not significantly related to a person’s general health

H1: ρ ≠ 0, Extroversion is significantly related to a person’s general health

Test statistic: t = r*((1-r*r)/(n-2))^0.5 = 0.6997*((1-0.6997*0.6997))/(6-2))^0.5 = 0.2499

Critical value (Using Excel function T.INV.2T(probability,df))) = T.INV.2T(0.05,4) = 2.7764

Since test statistic is less than critical value, we do not reject the null hypothesis.

So, extroversion is not significantly related to a person’s general health.